OpenStudy (anonymous):

i need help pls tell me answer if yall do you guys are the best pls Q: The graph below shows the relationship between pounds of dog food and total cost, in dollars, for the dog food. A coordinate graph is shown. The label on the x axis is Weight in pounds. The label on the y axis is Total cost in dollars. There is a line connecting the points 0, 0 and 10, 5. Select each statement about the graph that is true. A. Point (0, 0) shows the cost is $0 for 0 pound of dog food. B. Point (1, 0.5) shows the cost is $1.00 for 0.5 pound of dog food. C. Point (2, 1) shows that 2 pounds of

OpenStudy (anonymous):

pls anyone help me pls

OpenStudy (anonymous):

anyone pls

OpenStudy (skylyerbailey):

I am working on the answer right now...one second please :)

OpenStudy (skylyerbailey):

You're answers are A...C and E b is (1,0.5) so 0.5 = 1(1/2) which is true but it says "$1.00 for 0.5 pound" which is false since the x axis is weight c is (2,1) so 1 = 2(1/2) which is true, which is 2 pounds for $1. so c is correct d is (4,2) so 2 = 4(1/2) which is true, which is 4 pounds for $2, so d is not correct e is (10,5) so 5 = 10(1/2) which is true, which is 10 pounds for $5, so e is also correct

OpenStudy (anonymous):

thx bro

OpenStudy (skylyerbailey):

I hope that helps \(^_^)/

OpenStudy (skylyerbailey):

and I am a girl XD

OpenStudy (anonymous):

thx girl XD

OpenStudy (skylyerbailey):

You're welcome boy XD

OpenStudy (anonymous):

lol

OpenStudy (skylyerbailey):

I will fan you....LOL you seem fun

OpenStudy (anonymous):

thx i would love to have cool air lol

OpenStudy (skylyerbailey):

LOL hahaahahahaah you are funny!!

OpenStudy (anonymous):

thx ur nice

OpenStudy (anonymous):

it was correct =) thx ur awsome

OpenStudy (skylyerbailey):

LOL thanks! You are nice too!!!! I feel like I helped the right person today X'D

OpenStudy (anonymous):

and yes you did helped the right person today

OpenStudy (skylyerbailey):

YAYAAYYAYAYAYA I thought I was stupid!!!! I am glad you got it right.....and you are awe-inspiring as well!!!!!

OpenStudy (anonymous):

thx see you later

OpenStudy (skylyerbailey):

See you later!!!